OpenStudy (anonymous):

For which system of equations would elimination be the simplest technique to choose? A. y = 2x and x + 3y = 5 B. 4x + 3y = 6 and 3y = 5 C. x + y = 7 and x - y = 12 D. 3x - 2y = 8 and 4x - 5y = 20

hero (hero):

Hint: look for a pair of equations containing two equal yet opposite variables.

hero (hero):

2nd Hint: In the two equations below, the opposite variables + a and -a allow for simple variable elimination because +a - a = 0: x + a = 4 x - a = 2

hero (hero):

3rd Hint: 7x + y = 2 7x - y = 3 The two equations above contain opposite variables +y and - y. Since + y - y = 0, this allows for simple elimination.